# Coherent High-Harmonic Generation with Laser-Plasma Beams

### v1

@@ Abstract
While there are multiple ways of producing coherent radiation at a desired wavelength from an electron beam, LPA beams come with peculiar properties that make many conventional schemes technically challenging. For example, in a Self-Amplified Spontaneous Emission (SASE) [8] FEL scheme a combination of low emittance, high peak current, and low energy spread is required otherwise the FEL gain length may become too long for practical applications. Recently, the first breakthrough experiments have been done to demonstrate SASE at 27 nm with an LPA [5]. Yet, jitters and energy spread in the LPA beam might affect the quality of the photon radiation in terms of its bandwidth and wavelength stability. The energy spread and jitter can be drastically reduced by employing an active energy compression scheme [3, 9] that trades off the reduction in energy variation against a decompression in time. Consequently, achieving sufficiently low energy spreads may lead to insufficient peak beam currents, resulting in unacceptably long gain lengths. Synchrotron radiation can be enhanced by modulating the density of the beam [10], for example using a seed laser as demonstrated in [6]. But, operating at the seed wavelength naturally results in relatively long emission wavelength, which is limited by the seed laser.
Another approach, a low-gain FEL [11], takes advantage of optical cavity to build up photon radiation over multiple passages of the electron beam, thus increasing the gain length. Due to the relatively low, $<100$ Hz LPA repetition rates this scheme seems to be unfeasible.
@@
Alternatively, one may consider high-harmonic generation [12, 13, 14, 15, 16]. In this scheme, the electron beam is modulated by a laser pulse in the first undulator, passes through a dispersive chicane that convert energy modulation into longitudinal density modulation, and radiates in the last undulator. The key factor limiting the access to higher harmonics in this scheme is the uncorrelated energy spread within the bunch $\sigma_{E}$. Efficient generation of harmonics requires the energy modulation amplitude $A=\Delta E/\sigma_{E}\geq n$, making generation of high harmonics $n\gg 1$ particularly challenging.
@@
Active energy compression allows the LPA beam reaching extremely low levels of energy spread and jitter, down to $\sim 10^{-5}$, employing conventional radiofrequency (RF) accelerating cavities for active energy compression [3]. This scheme is envisioned for the PETRA IV plasma injector [1] to enable clean and efficient injection in the storage ring. Once filled, the storage ring needs to be topped up only once every few minutes to restore a small fraction ($\sim 1$) of the total charge lost due to Touschek and residual gas scattering  [^1] Contemporary and planned 4th generation light sources, such as PETRA IV feature large $\gtrsim 10$ h beam lifetimes.. This opens a window of opportunity for additional applications of an LPA injector that could take advantage of high quality electron beams, when they are not required for the storage ring. Thanks to their small energy spread after energy compression the LPA beams require a relatively small energy modulation to access high $\sim 100$-th harmonics of the seed wavelength, while a high power laser for the seeding is naturally present in the LPA setup. Thus, the LPA electron beams can produce stable, narrow-band coherent radiation at high harmonics of the driving laser pulse, achieving wavelengths as small as $\sim 10$ nm.
@@
The energy-compressed beam then enters the first undulator where it co-propagates with a seed laser pulse that is generated by splitting a small fraction of the drive pulse. The laser field imprints a modulation of the beam energy with a spatial period of the laser wavelength [Fig. 1(d)]. After passing though a dispersive section the energy modulation translates into microbunching with the periodicity of the modulation wavelength [Fig. 1(e)]. The resulting microbunches have a peak current far exceeding that of the initial bunch. And finally, the resulting train of short microbunches radiates in the second undulator at the high $n\gg 1$ harmonic of the seed laser wavelength $\lambda_{l}$.
@@
The power radiated at the $n$-th harmonic of the modulation wavelength through the last undulator is proportional to $a_{n}^{2}$ (see e.g. [20]). In order to provide an analytical estimation, we may consider the limit of a thin electron beam [21], such that the diffraction parameter $\sigma_{r}^{2}nk_{l}/L_{u}\ll 1$. In this regime, the radiation power is independent of the transverse beam size $\sigma_{r}$:
@@
Figure 5: Peak brilliance of the radiation calculated using the code Synchrad [26]. Blue (orange) curves show the spectrum of the simulation case I (II). Vertical dashed lines depict the harmonics of the seed laser.
@@
It has to be noted that high harmonics can also be generated by seeding the beam with two laser pulses in the echo-enabled harmonic generation (EEHG) scheme [28, 29]. While EEHG does not require a large modulation amplitude to generate high harmonics, compared to CHHG, we avoid this limitation thanks to the small $\sigma_{\delta}$ of the LPA beam. This allows us targeting the harmonics that are higher than what is typically achievable with the EEHG (e.g. $n\sim 30$ in Ref. [30]).
In conclusion, with the help of energy compression technique modern LPAs are approaching a point where they will be capable of delivering electron beams with low energy spread and jitter. We are proposing a scheme that takes advantage of this low energy spread to generate very high, $\sim 100$-th harmonics of the driver laser. Our estimates have shown that using a small fraction of the driver laser pulse to modulate the electron beam energy one can form a train of few thousands microbunches within the full width at half maximum of the initial electron bunch. This is possible thanks to the low, $10^{-5}$-level energy spread of the electron beam. The resulting short microbunches then allow accessing high harmonics of the modulator wavelength in the radiator undulator, producing $\sim 1$-ps-long bursts of radiation. The radiation wavelength can be easily tuned by selecting the desired harmonic of the seed $\lambda_{l}$, adjusting the radiator $K$ parameter or the beam energy. The distance between adjacent harmonics is about $\lambda_{l}/n^{2}\approx 0.1$ nm for the 80-th harmonic of 800 nm. Similar radiation power can be generated at each wavelength, because the bunching factor depends weakly on $n$ for high $n$’s (Fig. 3). Since the width of the resulting radiation is determined by the spectrum of the beam current, the long microbunch trains would generate extremely narrow spectral lines, as seen in our numerical simulation. Ultimately, one can aim at achieving the width as small as $\sim 10^{-5}$. In practice, the spectral width and phase noise of the seed laser pulse and the nonlinear energy chirp, remaining in the electron beam after the energy compression, will broaden the spectrum. Together these effects might worsen the relative width to $\sim 10^{-4}$ level. The narrow bandwidth might find applications in spectroscopy where a high spectral resolving power is required. The proposed scheme has a potential for scalability to higher beam energies, which can help mitigating detrimental effects of space charge and beam divergence. Since the radiation wavelength is defined by the harmonics of the seed wavelength, a higher beam energy might not result in a higher photon energy. Still, it may enable a wider range of undulator parameters such as a longer period and higher strength, allowing for additional tunability of the resulting radiation.

### v2

@@ Abstract
While there are multiple ways of producing coherent radiation at a desired wavelength from an electron beam, LPA beams come with peculiar properties that make many conventional schemes technically challenging. For example, in a Self-Amplified Spontaneous Emission (SASE) [8] FEL scheme a combination of low emittance, high peak current, and low energy spread is required, otherwise the FEL gain length may become too long for practical applications. Recently, the first breakthrough experiments have been done to demonstrate SASE at 27 nm with an LPA [5]. Yet, jitters and energy spread in the LPA beam might affect the quality of the photon radiation in terms of its bandwidth and wavelength stability. The energy spread and jitter can be drastically reduced by employing an active energy compression scheme [3, 9] that trades off the reduction in energy variation against a decompression in time. Consequently, achieving sufficiently low energy spreads may lead to insufficient peak beam currents, resulting in unacceptably long gain lengths. In principle, a low-gain FEL [10] could avoid this limitation by taking advantage of an optical cavity to build up radiation over multiple passages of the electron beam, thus decreasing the gain length. But due to the relatively low, $<100$ Hz LPA repetition rates this scheme seems to be unfeasible.
@@
Alternatively, synchrotron radiation can be enhanced by modulating the density of the beam [11], for example using a seed laser as demonstrated in [6]. However, the emission wavelength is typically limited by that of the seed laser to the visible spectrum. In order to significantly reduce the emission wavelength, one may consider high-harmonic generation [12, 13, 14, 15, 16]. In this scheme, the electron beam is modulated by a laser pulse in the first undulator, passes through a dispersive chicane that converts energy modulation into longitudinal density modulation, and radiates in the last undulator. The key factor limiting the access to higher harmonics in this scheme is the uncorrelated energy spread within the bunch $\sigma_{E}$. Efficient generation of harmonics requires an energy modulation amplitude $A=\Delta E/\sigma_{E}\geq n$, making generation of high harmonics $n\gg 1$ particularly challenging.
@@
Active energy compression allows the LPA beam reaching extremely low levels of energy spread and jitter, down to $\sim 10^{-5}$, employing conventional radiofrequency (RF) accelerating cavities for active energy compression [3]. This scheme is envisioned for the PETRA IV plasma injector [1] to enable clean and efficient injection in the storage ring. Once filled, the storage ring needs to be topped up only once every few minutes to restore a small fraction ($\sim 1$) of the total charge lost due to Touschek and residual gas scattering  [^1] Contemporary and planned 4th generation light sources, such as PETRA IV feature large $\gtrsim 10$ h beam lifetimes.. This opens a window of opportunity for additional applications of an LPA injector that could take advantage of high quality electron beams, when they are not required for the storage ring. Thanks to their small energy spread after energy compression the LPA beams require a relatively small energy modulation to access high $\sim 100$-th harmonics of the seed wavelength, while a high power laser for the seeding is naturally present in the LPA setup. Thus, the LPA electron beams can produce stable, narrow-band coherent radiation at high harmonics of the driving laser pulse, achieving wavelengths as small as 10 nm for an 800 nm seed.
@@
The energy-compressed beam then enters the first undulator where it co-propagates with a seed laser pulse that is generated by splitting a small fraction of the drive pulse. The laser field imprints a modulation of the beam energy with a spatial period of the laser wavelength [Fig. 1(d)]. After passing through a dispersive section the energy modulation translates into microbunching with the periodicity of the modulation wavelength [Fig. 1(e)]. The resulting microbunches have a peak current far exceeding that of the initial bunch. And finally, the resulting train of short microbunches radiates in the second undulator at the high $n\gg 1$ harmonic of the seed laser wavelength $\lambda_{l}$.
@@
The power radiated at the $n$-th harmonic of the modulation wavelength through the last undulator is proportional to $a_{n}^{2}$ (see e.g. Ref. [20]). In order to provide an analytical estimation, we may consider the limit of a thin electron beam [21], such that the diffraction parameter $\sigma_{r}^{2}nk_{l}/L_{u}\ll 1$. In this regime, the radiation power is independent of the transverse beam size $\sigma_{r}$:
@@
Figure 5: Peak brilliance of the radiation calculated using the code Synchrad [26]. The blue curve shows the spectrum of the simulation case I, $\lambda_{l}=800$ nm; the orange one shows the case II, $\lambda_{l}=270$ nm. Vertical dashed lines depict the harmonics of the seed laser.
@@
In conclusion, with the help of the energy compression technique modern LPAs are approaching a point where they will be capable of delivering electron beams with low energy spread and jitter. We are proposing a scheme that takes advantage of this low energy spread to generate very high, $\sim 100$-th harmonics of the driver laser. Our estimates have shown that using a small, $\sim 1\%$ fraction of the driver laser pulse to modulate the electron beam energy one can form a train of few thousands microbunches within the full width at half maximum of the initial electron bunch. This is possible thanks to the low, $10^{-5}$-level energy spread of the electron beam. The resulting short microbunches then allow accessing high harmonics of the modulator wavelength in the radiator undulator, producing $\sim 1$-ps-long bursts of radiation. The radiation wavelength can be easily tuned by selecting the desired harmonic of the seed $\lambda_{l}$, adjusting the radiator $K$ parameter or the beam energy. The distance between adjacent harmonics is about $\lambda_{l}/n^{2}\approx 0.1$ nm for the 80-th harmonic of 800 nm. Similar radiation power can be generated at each wavelength, because the bunching factor depends weakly on $n$ for high $n$’s (Fig. 3). Since the width of the resulting radiation is determined by the spectrum of the beam current, the long microbunch trains would generate extremely narrow spectral lines, as seen in our numerical simulation. Ultimately, one can aim at achieving the width as small as $\sim 10^{-5}$. In practice, the spectral width and phase noise of the seed laser pulse and the nonlinear energy chirp, remaining in the electron beam after the energy compression, will broaden the spectrum. Together these effects might worsen the relative width to $\sim 10^{-4}$ level. The narrow bandwidth might find applications in spectroscopy where a high spectral resolving power is required. The proposed scheme has a potential for scalability to higher beam energies, which can help mitigating detrimental effects of space charge and beam divergence. Since the radiation wavelength is defined by the harmonics of the seed wavelength, a higher beam energy may not always result in a higher photon energy. Still, it may enable a wider range of undulator parameters such as a longer period and higher strength, allowing for additional tunability of the resulting radiation.